1 kg of steam initially dry saturated at 1.1 MPa expands in a cylinder as per pv1.13 = Constant. The pressure at the end of expansion is 0.1 MPa. Determine

(a) The final volume

(b) Final dryness fraction

(c) Workdone and

(d) Heat transferred


1
Expert's answer
2021-09-03T01:59:35-0400

To find; The final volume, Final dryness fraction,Workdone and heat transferred.


A = πD2/4 = π x (0.04)2/4 = 0.001257 m2

p=200bar θ=400oC V=mv=0.00995m3/s

u = V/A = 0.00995/0.001257 = 7.91 m/s

KE = mu2/2 = 1 x 7.912/2 = 31.3 W

h = 2819 kJ/kg 


Take cp as 1005 J/kg

K. Φ+P=∆H/s (ignorePEandKE)

∆H/s=mcp

∆T=0.7x1.005(120–20)=70.35kW

5 + P = 70.35 P = 65.35 kW